Solve for x and y in the following system of equations:
2x + 3y = 5
3x + 4y = 6

Answer by G Nana(3)

2x+3y=5
3x+4y=6
You multiply the first equation by -4 and the second equation by 3 to get the value of x.
So, you will
-8x-12y=-20
9x+12y=18
-12y+12y =0
you will get
-8x=-20
9x=18
by substitution-8x+9x=x and -20+18=-2 then x= -2
You apply the value of x in first equation
2(-2)+3y= 5
-4+3y= 5
3y= 5+4
3y= 9
y= 9/3= 3
So the answer is x=-2 and y= 3